1. Given a normally distributed population with mean of 100 and a standard deviation of 20, find P(96 ≤ x-bar ≤ 108). Sample size is 16.
Answer by stanbon(75887)

Given a normally distributed population with mean of 100 and a standard deviation of 20, find P(96 ≤ x-bar ≤ 108). Sample size is 16.
-----------------
z(96) = (96-100)/(20/sqrt(16)) = -4/5
---
z(108) = (108-100)/(20/sqrt(16)) = 8/5
-------
P(96<= x-bar <=108) = P(-4/5<= z <=8/5) = 0.7334
-----------------
Cheers,
Stan H.
--------------